OpenStudy (anonymous):

What is the value of n in the equation 8n + 6 = 10? Please enter your answer as a fraction.

OpenStudy (doc.brown):

\[8n+6=10\]\[8n=10-6\]\[\frac{8n}{8}=\frac{10-6}{8}\]